OpenStudy (anonymous):

simplify by factoring 3sqrt243x^8

OpenStudy (anonymous):

I dont' understand your expression; is the x^8 in the radicand?

OpenStudy (anonymous):

Do you have \[\sqrt[3]{243x^8}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then you have\[\sqrt[3]{243x^8}=\sqrt[3]{3^5x^8}=3x^2\sqrt[3]{9x^2}\]

OpenStudy (anonymous):

:-) thanks

OpenStudy (anonymous):

my pleasure

OpenStudy (anonymous):

can we do more

OpenStudy (anonymous):

ask away